MFA/”Emerald Necklace” Conservancy Trip – Saturday, April 16

Schedule 8:45 am – Busses leave from Hogan 3 (upper entrance opposite parking lot) 10 am – Arrival at Forsyth Street and MFA 10 am to 12 noon – MFA (Mayan collection in Art of Americas Wing, and anything else you want or have time to see) 12 noon – Lunch (provided) at Back Bay Fens (or Northeastern Marino Center on Huntington Avenue if weather does not cooperate) 1 pm - 2:30 pm – Walking tour of Back Bay Fens Park in “Emerald Necklace” roughly 2:45 – Busses leave Forsyth Street for Fanueil Hall roughly 3:30 – 5:00 pm – Free time at Fanueil Hall (shopping, etc.) 5:00 pm – Busses leave Fanueil Hall for return to Holy Cross roughly 6:15 – Back to campus/

Suggestions

1. Bring paper, writing implements to take notes while you are in the MFA, and a copy of the writing assignment.

2. Bring a camera

3. Exchange cell phone numbers with a “buddy” so you can get in contact if you get separated from the group.